40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Dec 1 07:30:29 2014, Steve, Update, SUS, ETMX damping restored 
    Reply  Mon Dec 1 16:34:20 2014, Jenne, Update, General, IFO wake-up 
       Reply  Mon Dec 1 17:19:22 2014, Jenne, Update, General, PMC align 
Message ID: 10742     Entry time: Mon Dec 1 17:19:22 2014     In reply to: 10740
Author: Jenne 
Type: Update 
Category: General 
Subject: PMC align 

[Diego, Jenne]

Tweaked up the input alignment to the PMC.  Now we're at 0.785.

ELOG V3.1.3-